OpenStudy (anonymous):

if i have (y-39)(2y+7)(y+12)=0 what do i do with the 2y+7?

OpenStudy (saifoo.khan):

what you have to find?

OpenStudy (diyadiya):

" 0 "

OpenStudy (anonymous):

You set each group (the numbers in parenthesis) to 0 so like y-39=0, 2y+7= 0, and y+12=0 . Find the 3 values of y and that is your answer

OpenStudy (anonymous):

i have to solve using the zero factor property

OpenStudy (saifoo.khan):

Alright, then: (y-39)(2y+7)(y+12)=0 y = 39 y = -7/2 y = -12